OpenStudy (kj4uts):

Given four different data sets, a student calculates the quadratic regression as well as the R^2 value for each using a TI-83...?

OpenStudy (kj4uts):

Given four different data sets, a student calculates the quadratic regression as well as the R^2 value for each using a TI-83. Which of these result models its respective data the best? Set A: a = 1.23 b = -3.27 c = 7.83 R^2 = 0.87394 Set B: a = -3.745 b = 9.524 c = 12.105 R^2=0.1420 Set C: a = 142.3629 b = 204.2284 c = -428.2079 R^2 = 0.9328 Set D: a = 35.29 b = 29.94 c = -10.23 R^2 = 0.0139 Please explain. Thank you!

OpenStudy (dumbcow):

R^2 indicates how well the regression equation fits the data R^2 = 1 means it perfectly fits the data R^2 = 0 means it doesn't fit the data at all